OpenStudy (anonymous):

(7x)^09^79^x=9^2 solve for x

OpenStudy (anonymous):

is that 7x raised to all those exponents?

myininaya (myininaya):

\[(7x)^09^79^x=9^2 \]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ah ok then.

myininaya (myininaya):

Assuming x does not equal 0 what is \[(7x)^0\]?

OpenStudy (anonymous):

idk.

myininaya (myininaya):

Do you what \[5^0=?\]

myininaya (myininaya):

know*

OpenStudy (anonymous):

yea

myininaya (myininaya):

what does it equal?

myininaya (myininaya):

\[(anything)^0=1 , \text{ where } anything \neq 0 \]

myininaya (myininaya):

So \[5^0=1\] And \[(7x)^0=1, \text{ assuming } x \neq 0\]

myininaya (myininaya):

\[(7x)^09^{7}9^x=9^2\] \[1 \cdot 9^7 9^x=9^2\] \[9^7 9^x=9^2\] Do you know what to do now?

OpenStudy (anonymous):

:D yess. thanks